Construct a confidence interval for

p1−p2 at the given level of confidence.

x1=365​,

n1=503​,

x2=447​,

n2=558​,

95​% confidence
Answer #1

Solution , Given that X = 365 ni = 503 X2=447 n2 = 558 Ô - XL365 -0.7256 503 I-pi= 1-0.7256 = 0.2744 ệz = x2 - 441 -0.8011 55